Calculate the molality of the solution obtained by dissolving 11.7g NaCl in 500g water

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
Molality = moles(solute)/ kg(solvent)
 = (11.7/58.4)/ 0.5
 = 0.2/0.5 = 0.4
 
Regards
Arun (askIITians forum expert)